# The shadow and observational images of the non-singular rotating black holes in loop quantum gravity

### v1

@@ 1 Introduction
Since 2019, the Event Horizon Telescope (EHT) international collaboration team has obtained images of the black hole at the center of the radio galaxy M87* and the Milky Way’s SgrA* [1, 2]. Both images share similarities, featuring a relatively dark central region and a brighter outer ring structure. These two regions are referred to as the black hole shadow and the photon ring, respectively. In fact, light near a black hole is absorbed by the black hole and cannot reach the observer, resulting in a dark region against the background sky when observing the black hole [3]. This dark region is the black hole shadow. Strictly speaking, due to the deflection effects of the gravitational field on light, the dark region within the critical curve is defined as the black hole shadow, while light approaching the critical curve from infinity will asymptotically approach the boundary photon orbit [4]. The advent of black hole images marks a significant milestone in the history of black hole research and holds substantial scientific significance. These images not only provide valuable information about the accretion processes, radiation mechanisms, and jet mechanisms near black holes, but also extract the spacetime characteristics of black holes. For a long time, the black hole shadow and its observable effects have been one of the focal points of research. For example, the shadow of a Schwarzschild black hole appears as a black disk from any angle, while an observer in the equatorial plane will notice that the shadow of a Kerr black hole changes shape with the rotation parameter. As the rotation parameter increases, the black disk gradually evolves into a “D” shape [5]. Later on, detailed studies have been conducted on the shadows of rotating EMDA black hole [6], non-commutative black holes [7], phantom black holes [8], Konoplya-Zhidenko black holes [9], and black hole with massive vector fields [10], revealing the double shadows, cuspy shadow structures and other interesting phenomenon [11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. To capture a more detailed view of black hole shadows and the light distortion behavior around them, based on a four-color celestial sphere light source model, one has numerically studied Einstein rings caused by gravitational lensing effects, as well as images of black hole shadows and the space-dragging effect caused by rotation [21]. This method was later extended to black hole models in various gravitational theories, used for numerically simulating black hole shadows when the photon motion system is integrable [22]. In fact, black holes in the universe are always surrounded by various accreting materials [23]. In view of this, when a simple static optically thin spherical accretion surrounds a Schwarzschild black hole, Narayan et al. discovered in 2019 that the black hole shadow and photon sphere reflect the geometry of spacetime, and their sizes are almost unaffected by the accreting material [24]. In fact, there is another typical disk-shaped accreting matter exists in the universe, namely the accretion disk. And, the astronomical observations indicate that the accretion disk indeed exists around the supermassive black hole SgrA* at the center of the Milky Way. In 2019, Wald et.al discovered that when a Schwarzschild black hole is surrounded by the accretion disk (the geometrically and optically thin disk), the observer located at the North pole would find the black hole’s shadow as a dark disk encircled by a bright ring. This bright ring is composed of direct emissions, a lensing ring, and a photon ring [25]. Later on, this idea was further extended to other static spherically symmetric black holes containing matter fields, or to black holes or wormhole spacetimes in other gravitational theories [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. It was proposed that the structure of the black hole shadow and the associated rings could be used to distinguish between different black hole models in various gravitational theories. Considering the strong magnetic field around black holes, in 2022, this work [25] was extended for the first time to the case of rotating Kerr black holes, analyzing the effects of rotation, magnetic fields, and the observer’s angle on the multi-level images of black holes [39]. In [39], it was assumed that the accretion disk around the Kerr black hole is composed of plasma fluid and is located in the equatorial plane of black hole. The fluid moves along Keplerian orbits outside the innermost stable circular orbit (ISCO), while inside the ISCO, it moves along simplified precessing geodesic trajectories with the energy and angular momentum of the ISCO. Of course, in recent years, some progresses have been made in the study of black hole shadows and images, particularly in exploratory research on hotspot images [40, 41], polarized images [42, 43, 44], jet images [45], boson star images [46], accretion disk models [47], and so on. At present, by considering the thin disk models [39], it should be noted that the rotating black hole images in other gravity models and matter fields remains unknown. Therefore, continuing to study black hole images under this model would be a highly significant endeavor, as it may provide an effective tool for distinguishing different gravitational models and exploring quantum effects.
@@
On the other hand, black hole singularitity is the point in spacetime where the curvature becomes infinite, at which the mass of matter is compressed infinitely, with the density approaching to infinity and the volume approaching zero. And, Penrose have proven that all matter within a black hole should collapse into the singularity [48]. The formation of singularities is almost inevitable within the framework of classical general relativity, making it impossible to predict the evolution of physical phenomena in spacetime effectively. Currently, it is widely believed that the existence of spacetime singularities reflects the incompleteness of general relativity, and that resolving the singularity problem likely requires a theory of quantum gravity(QG). To date, a complete theory describing QG has yet to be found. This has prompted physicists to focus on constructing nonsingular black holes within classical general relativity. Typically, one can seek nonsingular solutions to the field equations to address the singularity problem, at the cost of violating the strong energy condition. These solutions lack intrinsic singularities at the coordinate origin, and event horizons can still exist. Such black holes are referred to as regular black holes or non-singular black holes. In 1968, Bardeen proposed the first nonsingular black hole, namely Bardeen black hole, which satisfies the weak energy condition [49]. In 2000, it was shown that the physical motivation for the absence of spacetime singularities in this solution is that the Bardeen black hole is a spherically symmetric, static solution derived from coupling a nonlinear electrodynamics model to Einstein’s gravitational field equations [50]. Subsequently, many similar spherically symmetric regular black holes have been constructed, such as the Hayward black hole, the Ayón-Beato-García black hole [51], the Berej-Matyjasek-Trynieki-Wornowicz black hole [52], and the Simpson-Visser black hole [53], among others [54, 55, 56, 57, 58, 59, 60, 61, 62]. Since regular black holes are considered manifestations of QG effects, various properties of regular black holes have been extensively studied, such as quasinormal modes [63], greybody factors [64], superradiance [65], Joule-Thomson expansion [66], P-V criticality [67], dynamical stability [68], gravitational lensing [69], shadow profiles [70], and spherical accretion shadow appearances [71], among others. By exploring these properties, it is possible for us to capture effects related to QG. Currently, the study of regular black holes has become one of the focal points and hot topics in contemporary research.
@@
The loop quantum gravity(LQG), as one of the main candidates for QG theory, is a non-perturbative approach to QG that introduces new dynamical variables in a connection dynamics framework, which has drawn a great deal of attention [72, 73, 74, 75, 76, 77, 78, 79, 80, 81]. This theory not only has a well-defined Hilbert space for LQG dynamics but also naturally predicts a discrete structure of spacetime geometry at the Planck scale. More importantly, the non-rotating LQG black hole solutions obtained in this gravitational theory exhibits a non-singular geometry that we expect due to the existence of transition surface. By using the Newman-Janis-Algorithm(NJA) that is a solution-generating method [82], one has constructed the rotating black hole solutions from the Schwarzschild black hole [83]. This solution does not depend on the specific details of the seed metric used, thereby effectively capturing some universal properties of rotating LQG black holes. At present, the dynamical behavior of a scalar field near this black hole has been investigated [84]. However, it remains unclear whether the non-singular effects of this black hole (or the influence of the transition surface) and quantum effects have observable astronomical consequences. Therefore, based on the current progress in black hole images, it is a highly important research for studying the shadow, inner shadow, celestial source images, and thin disk images of non-rotating LQG black holes by using the thin disk accretion model. In this context, this paper will focus on the astronomical observable effects of non-rotating LQG black holes, providing potential references for distinguishing non-singular black holes and analyzing the quantum effects of black holes.
@@ 2 Review of the rotating LQG black holes
In this section, we briefly review the non-singular rotating black holes in loop QG. Because of the complexity of Einstein’s field equations in the rotating case, the Newman-Janis algorithm(NJA) method is widely accepted during the construction of the rotating black holes. The solution of the rotating LQG black holes we focused on in this paper is obtained by this method. In general, the seed metric in a general static and spherically symmetric space time in 4-dimensional case always reads
@@
Based on the NJA, by introducing the advanced null coordinates $(u,t,\theta,\phi)$, where $du=dt-\frac{dr}{\sqrt{f(r)g(r)}}$, one can use a null tetrad ($Z_{a}^{\mu}=(l^{\mu},n^{\nu},m^{\mu},\bar{m}^{\mu})$) to express the inverse of the seed metric,
@@
In Eq.(2), $\bar{m}^{\mu}$ is the complex conjugate of ${m}^{\mu}$. The null tetrad in the advanced null coordinates can be expressed as following,
$l^{\mu}=\delta_{r}^{\mu},\quad n^{\mu}=\sqrt{\frac{f(r)}{g(r)}}\delta_{\mu}^{\mu}-\frac{f(r)}{2}\delta_{r}^{\mu},\quad m^{\mu}=\frac{1}{\sqrt{2h(r)}}\left(\delta_{\theta}^{\mu}+\frac{i}{\sin\theta}\delta_{\phi}^{\mu}\right)$ (3)
@@
$\delta_{r}^{\mu}\rightarrow\delta_{r}^{\mu},\quad\delta_{u}^{\mu}\rightarrow\delta_{u}^{\mu},\quad\delta_{\theta}^{\mu}\rightarrow\delta_{\theta}^{\mu}+ia\sin\theta(\delta_{u}^{\mu}-\delta_{r}^{\mu}),\quad\delta_{\phi}^{\mu}\rightarrow\delta_{\phi}^{\mu}$ (6)
@@
$ds^{2}$ $=-Adu^{2}-2\sqrt{\frac{A}{B}}dudr-2a\sin^{2}\theta\left(\sqrt{\frac{A}{B}}-A\right)dud\phi+2a\sin^{2}\theta\sqrt{\frac{A}{B}}drd\phi+\Psi d\theta^{2}$ (7)
@@
With the help of the coordinate transformation and some constraints in [85], we can return to the Boyer-Lindquist coordinates, and can only considered one off-diagonal term $g_{t\phi}$ in above metric. Finally, when one reconsidered the seed metric in LQG that described in [86, 83], the rotating black holes with the Kerr-like form can be expressed as,
@@
In fact, $\Psi(r,\theta,a)$ should be determined by some especial physical interpretations. However, here when the source is interpreted as an imperfect fluid rotating about the $z$ axis, $\Psi(r,\theta,a)$ is obtained, which is $\Psi(r,\theta,a)=\Sigma=\mathcal{H}(r)+a^{2}\cos^{2}{\theta}$. The quantum parameter $\lambda$ originates from holonomy modifications, which represents the effects of QG. For the rotating regular black hole (9), the horizons reads
@@ 3 Shadows of the rotating LQG black holes
where $H$ is the canonical Hamiltonian, $p_{\mu}$ is the four momentum. By considering the symmetries of the spacetime and the associated Killing vectors, there are two conserved quantities along a geodesic, namely, $E$ and $L$, which represent the energy and angular momentum in the direction of the axis of symmetry. They are,
@@
The symbol $\dot{t}$ represents the derivative of $\tau$, which is the affine parameter. For null geodesics, i.e., $m=0$, the equations of motion for photons propagating in the spacetime (9) can be obtained by solving the Hamilton-Jacobi equation. Further combined with Eq.(15), the equations of motion for photons can be expressed as the following four first-order differential equations,
@@
$R=\left[a^{2}-a\xi+\mathcal{H}\right]^{2}E^{2}-\Delta E^{2}\left(a^{2}-2a\xi+\eta+\xi^{2}\right)=0$ (20)
@@
$\xi=\frac{a^{2}\Delta^{\prime}-2\Delta\mathcal{H}^{\prime}+\mathcal{H}\Delta^{\prime}}{a\Delta^{\prime}}|_{r=r_{p}}$ (22)
@@
Due to the non-negativity of $\Theta$ in Eq.(16), it immediately presents the condition for the photon region,
@@
$e_{3}=-e_{\phi)}=\left(0,0,0,\frac{-1}{\sqrt{g_{\phi\phi}}}\right),$ (28)
@@
The observer is located at ($r_{O},\theta_{O}$) in the coordinates ($t,r,\theta,\phi$), which corresponding to the point $O$ in Fig.2. The red line with arrow originated from $O$ represents the propagated direction of the light ray. And, the vector $\overrightarrow{OP}$ denotes the tangent vector of the null geodesic at $O$. To further obtain the shadow and image of black hole, we take the $\overrightarrow{OP}$ as a radius and $O^{\prime}$ as the centre to plot a three-dimensional sphere, and place the origin of the tetrad (25) at $O^{\prime}$. The line $O^{\prime}S$ passed through the point $O$, as the diameter of the sphere, is on a line with $e_{1}$. The blue rectangular plane at $O^{\prime}$ is straight down, which is the screen for presenting the shadow and image of black hole. By regarding the point $S$ as the reference point, the vector $\overrightarrow{OP}$ is projected to be the vector ${O^{\prime}P^{\prime}}$ on the screen. To determine the photon’s position from the observer’s perspective, it is useful to introduce celestial coordinates($\alpha,\beta$). For the plane $O^{\prime}PS$ with the green boundary of the sphere, the first celestial coordinate, $\alpha$, is defined as the angle between $OO^{\prime}$ and $OP$. For the brown vertical plane of the sphere, it intersects with the front side of the plane $O^{\prime}PS$ at point $Q$. The other celestial coordinate, $\beta$, is defined as the angle between $OQ$ and $OP$.
For the null geodesic, i.e., $s(\tau)=(t(\tau),r(\tau),\theta(\tau),\phi(\tau))$, its tangent should be a linear combination of $e_{i}$ere $i=0,1,2,3.$, which form can be expressed as
@@
where, the negative sign ensures that the tangent vector is oriented towards the past, and the symbol $\cdot$ denotes the partial derivative with respect to the affine parameter $\tau$. And, we note that the path of the photon does not depend on its energy. Therefore, we set the energy of photon in the camera’s frame, $E$, to be unity, i.e., $E_{camera}=1=|\overrightarrow{OP}|\cdot\chi=-\frac{E}{\sqrt{g_{tt}}}|_{(r_{O},\theta_{O})}$. In addition, the four-momentum of photons in the ZAMO’s frame can be rewrote as $p_{(\mu)}=p_{\nu}e^{\nu}_{(\mu)}$, where $e^{\nu}_{(\mu)}$ is given by Eqs.(25-28). Since the values of the 4-momentum of a photon can be obtained by Eqs.(16-19), we can easy extract the momentum of photons $p_{(\mu)}$. Next, by following the method in [87], the connection between the celestial coordinates and the four-momentum of photons $p_{(\mu)}$ is represented as $\cos\alpha={p^{(1)}}/{p^{(0)}},\tan\beta={p^{(3)}}/{p^{(2)}}.$ Specifically, this two celestial coordinates for the rotating LQG black holes are,
@@
where, $\mathcal{N}=\mathcal{A}-a^{2}\Delta\cos^{2}\theta$, $\mathcal{M}=\mathcal{A}a^{2}\sin^{2}\theta(\mathcal{H}-\mathcal{F}\mathcal{H})^{2}$. With the help of Eq.(32), we introduce the Cartesian coordinates system ($x,y$) on the screen, it yields,
@@
In Fig.3, it shows the influence of $\lambda$ on the black hole shadow for different values of $a$. Clearly, one can see that the parameter $\lambda$ always decreases the size of shadow for both $a=0.1$ and $a=0.998$, and seems to have no effects on the deformation of shadow. According to the Fig.1, we know that the acceptable range of values for $\lambda$ is bigger and bigger with the decrease of the rotating parameter $a$. So, when $a=0.1$, the broad range of $\lambda$ leads to substantial variations in the shadow’s size. Conversely, with $a=0.998$, the narrower range of $b$ results in a smaller changes in the shadow size.
(a) $(a):r_{o}=100,\theta_{o}=\pi/2$
(b) $(b):a=0.998,r_{o}=100$
(c) $(c):a=0.998,\theta_{o}=\pi/2$
Figure 4: The shadow versus parameters $a,r_{o},\theta_{o}$.
@@
Next, we continue to focus on the deviation from the circularity ($\delta_{s}$) and the size ($R_{s}$) of the shadow cast by the rotating black holes, which defined by Hioki and Maeda [88].
@@
In Fig.5, the five reference points ($\alpha_{t},\beta_{t}$), ($\alpha_{b},\beta_{b}$),($\alpha_{r},0$), ($\alpha_{p},0$) and ($\alpha_{p^{\prime}},0$) correspond to the top, bottom, rightmost, and leftmost points of the shadow, leftmost point of the reference circle, respectively. The deviation from the circularity ($\delta_{s}$) and the size ($R_{s}$) are defined as,
@@
In Tables 1 and 2, we intuitively show the deviation from the circularity ($\delta_{s}$) and the size ($R_{s}$) of the shadow cast by the rotating regular black holes.

### v2

@@ 1 Introduction
Since 2019, the Event Horizon Telescope (EHT) international collaboration team has obtained images of the black hole at the center of the radio galaxy M87* and the Milky Way’s SgrA* [1, 2]. Both images share similarities, featuring a relatively dark central region and a brighter outer ring structure. These two regions are referred to as the black hole shadow and the photon ring, respectively. In fact, light near a black hole is absorbed by the black hole and cannot reach the observer, resulting in a dark region against the background sky when observing the black hole [3]. This dark region is the black hole shadow. Strictly speaking, due to the deflection effects of the gravitational field on light, the dark region within the critical curve is defined as the black hole shadow, while light approaching the critical curve from infinity will asymptotically approach the boundary photon orbit [4]. The advent of black hole images marks a significant milestone in the history of black hole research and holds substantial scientific significance. These images not only provide valuable information about the accretion processes, radiation mechanisms, and jet mechanisms near black holes, but also extract the spacetime characteristics of black holes. For a long time, the black hole shadow and its observable effects have been one of the focal points of research. For example, the shadow of a Schwarzschild black hole appears as a black disk from any angle, while an observer in the equatorial plane will notice that the shadow of a Kerr black hole changes shape with the rotation parameter. As the rotation parameter increases, the black disk gradually evolves into a “D” shape [5]. Later on, detailed studies have been conducted on the shadows of rotating EMDA black hole [6], non-commutative black holes [7], phantom black holes [8], Konoplya-Zhidenko black holes [9], and black hole with massive vector fields [10], revealing the double shadows, cuspy shadow structures and other interesting phenomenon [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Based on black hole shadow, one has explored various aspects of black hole, i.e., the quasinormal modes and parameter constraints of black holes, among other aspects, which offer deep insights into our understanding and comprehension of gravity [22, 23, 24, 25, 26, 27, 28, 35, 29, 30, 31, 32, 33, 34, 36, 37, 38]. To capture a more detailed view of black hole shadows and the light distortion behavior around them, based on a four-color celestial sphere light source model, one has numerically studied Einstein rings caused by gravitational lensing effects, as well as images of black hole shadows and the space-dragging effect caused by rotation [39]. This method was later extended to black hole models in various gravitational theories, used for numerically simulating black hole shadows when the photon motion system is integrable [40]. In fact, black holes in the universe are always surrounded by various accreting materials [41]. In view of this, when a simple static optically thin spherical accretion surrounds a Schwarzschild black hole, Narayan et al. discovered in 2019 that the black hole shadow and photon sphere reflect the geometry of spacetime, and their sizes are almost unaffected by the accreting material [42]. In fact, there is another typical disk-shaped accreting matter exists in the universe, namely the accretion disk. And, the astronomical observations indicate that the accretion disk indeed exists around the supermassive black hole SgrA* at the center of the Milky Way. In 2019, Wald et.al discovered that when a Schwarzschild black hole is surrounded by the accretion disk (the geometrically and optically thin disk), the observer located at the North pole would find the black hole’s shadow as a dark disk encircled by a bright ring. This bright ring is composed of direct emissions, a lensing ring, and a photon ring [43]. Later on, this idea was further extended to other static spherically symmetric black holes containing matter fields and modified grvities, or to boson stars or wormhole spacetimes, and so so on [44, 45, 46, 47, 48, 49, 57, 58, 59, 60, 50, 51, 52, 53, 54, 55, 56]. It was proposed that the structure of the black hole shadow and the associated rings could be used to distinguish between different black hole models in various gravitational theories. Considering the strong magnetic field around black holes, in 2022, this work [43] was extended for the first time to the case of rotating Kerr black holes, analyzing the effects of rotation, magnetic fields, and the observer’s angle on the multi-level images of black holes [61]. In [61], it was assumed that the accretion disk around the Kerr black hole is composed of plasma fluid and is located in the equatorial plane of black hole. The fluid moves along Keplerian orbits outside the innermost stable circular orbit (ISCO), while inside the ISCO, it moves along simplified precessing geodesic trajectories with the energy and angular momentum of the ISCO. Of course, in recent years, some progresses have been made in the study of black hole shadows and images, particularly in exploratory research on hotspot images [62, 63], polarized images [64, 65, 66], jet images [67], boson star images [68], accretion disk models [69], and so on. At present, by considering the thin disk models [61], it should be noted that the rotating black hole images in other gravity models and matter fields remains unknown. Therefore, continuing to study black hole images under this model would be a highly significant endeavor, as it may provide an effective tool for distinguishing different gravitational models and exploring quantum effects.
@@
On the other hand, black hole singularitity is the point in spacetime where the curvature becomes infinite, at which the mass of matter is compressed infinitely, with the density approaching to infinity and the volume approaching zero. And, Penrose have proven that all matter within a black hole should collapse into the singularity [70]. The formation of singularities is almost inevitable within the framework of classical general relativity, making it impossible to predict the evolution of physical phenomena in spacetime effectively. Currently, it is widely believed that the existence of spacetime singularities reflects the incompleteness of general relativity, and that resolving the singularity problem likely requires a theory of quantum gravity(QG). To date, a complete theory describing QG has yet to be found. This has prompted physicists to focus on constructing nonsingular black holes within classical general relativity. Typically, one can seek nonsingular solutions to the field equations to address the singularity problem, at the cost of violating the strong energy condition. These solutions lack intrinsic singularities at the coordinate origin, and event horizons can still exist. Such black holes are referred to as non-singular black holes or non-singular black holes. In 1968, Bardeen proposed the first nonsingular black hole, namely Bardeen black hole, which satisfies the weak energy condition [71]. In 2000, it was shown that the physical motivation for the absence of spacetime singularities in this solution is that the Bardeen black hole is a spherically symmetric, static solution derived from coupling a nonlinear electrodynamics model to Einstein’s gravitational field equations [72]. Subsequently, many similar spherically symmetric non-singular black holes have been constructed, such as the Hayward black hole, the Ayón-Beato-García black hole [73], the Berej-Matyjasek-Trynieki-Wornowicz black hole [74], and the Simpson-Visser black hole [75], among others [76, 77, 78, 79, 80, 81, 82, 83, 84]. Since non-singular black holes are considered manifestations of QG effects, various properties of non-singular black holes have been extensively studied, such as quasinormal modes [85], greybody factors [86], superradiance [87], Joule-Thomson expansion [88], P-V criticality [89], dynamical stability [90], gravitational lensing [91], shadow profiles [92], and spherical accretion shadow appearances [93], among others. By exploring these properties, it is possible for us to capture effects related to QG. Currently, the study of non-singular black holes has become one of the focal points and hot topics in contemporary research.
@@
The loop quantum gravity(LQG), as one of the main candidates for QG theory, is a non-perturbative approach to QG that introduces new dynamical variables in a connection dynamics framework, which has drawn a great deal of attention [94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104]. This theory not only has a well-defined Hilbert space for LQG dynamics but also naturally predicts a discrete structure of spacetime geometry at the Planck scale. More importantly, the non-rotating LQG black hole solutions obtained in this gravitational theory exhibits a non-singular geometry that we expect due to the existence of transition surface. By using the Newman-Janis-Algorithm(NJA) that is a solution-generating method [105], one has constructed the rotating black hole solutions from the Schwarzschild black hole [106]. This solution does not depend on the specific details of the seed metric used, thereby effectively capturing some universal properties of rotating LQG black holes. At present, the dynamical behavior of a scalar field near this black hole has been investigated [107]. However, it remains unclear whether the non-singular effects of this black hole (or the influence of the transition surface) and quantum effects have observable astronomical consequences. Therefore, based on the current progress in black hole images, it is a highly important research for studying the shadow, inner shadow, celestial source images, and thin disk images of non-rotating LQG black holes by using the thin disk accretion model. In this context, this paper will focus on the astronomical observable effects of non-rotating LQG black holes, providing potential references for distinguishing non-singular black holes and analyzing the quantum effects of black holes.
@@ 2 Review of the rotating LQG black holes
In this section, we briefly review the non-singular rotating black holes in LQG. Because of the complexity of Einstein’s field equations in the rotating case, the Newman-Janis algorithm(NJA) method is widely accepted during the construction of the rotating black holes. The solution of the rotating LQG black holes we focused on in this paper is obtained by this method. In general, the seed metric in a general static and spherically symmetric space time in 4-dimensional case always reads
@@
Based on the NJA, by introducing the advanced null coordinates $(u,t,\theta,\phi)$ where $du=dt-\frac{dr}{\sqrt{f(r)g(r)}}$, one can use a null tetrad ($Z_{a}^{\mu}=(l^{\mu},n^{\nu},m^{\mu},\bar{m}^{\mu})$) to express the inverse of the seed metric
@@
In Eq.(2), $\bar{m}^{\mu}$ is the complex conjugate of ${m}^{\mu}$. The null tetrad in the advanced null coordinates can be expressed as following
$l^{\mu}=\delta_{r}^{\mu},\quad n^{\mu}=\sqrt{\frac{f(r)}{g(r)}}\delta_{\mu}^{\mu}-\frac{f(r)}{2}\delta_{r}^{\mu},\quad m^{\mu}=\frac{1}{\sqrt{2h(r)}}\left(\delta_{\theta}^{\mu}+\frac{i}{\sin\theta}\delta_{\phi}^{\mu}\right),$ (3)
@@
$\delta_{r}^{\mu}\rightarrow\delta_{r}^{\mu},\quad\delta_{u}^{\mu}\rightarrow\delta_{u}^{\mu},\quad\delta_{\theta}^{\mu}\rightarrow\delta_{\theta}^{\mu}+ia\sin\theta(\delta_{u}^{\mu}-\delta_{r}^{\mu}),\quad\delta_{\phi}^{\mu}\rightarrow\delta_{\phi}^{\mu}.$ (6)
@@
$ds^{2}$ $=-Adu^{2}-2\sqrt{\frac{A}{B}}dudr-2a\sin^{2}\theta\left(\sqrt{\frac{A}{B}}-A\right)dud\phi+2a\sin^{2}\theta\sqrt{\frac{A}{B}}drd\phi+\Psi d\theta^{2}$
@@
With the help of the coordinate transformation and some constraints in [108], we can return to the Boyer-Lindquist coordinates, and can only considered one off-diagonal term $g_{t\phi}$ in above metric. Finally, when one reconsidered the seed metric in LQG that described in [109, 106], the rotating black holes with the Kerr-like form can be expressed as
@@
In fact, $\Psi(r,\theta,a)$ should be determined by some especial physical interpretations. However, here when the source is interpreted as an imperfect fluid rotating about the $z$ axis, $\Psi(r,\theta,a)$ is obtained, which is $\Psi(r,\theta,a)=\Sigma=\mathcal{H}(r)+a^{2}\cos^{2}{\theta}$. The quantum parameter $\lambda$ originates from holonomy modifications, which represents the effects of QG. For the non-singular rotating black hole (8), the horizons reads
@@ 3 Shadows of the rotating LQG black holes
where $H$ is the canonical Hamiltonian, $p_{\mu}$ is the four momentum. By considering the symmetries of the spacetime and the associated Killing vectors, there are two conserved quantities along a geodesic, namely, $E$ and $L$, which represent the energy and angular momentum in the direction of the axis of symmetry. They are
@@
The symbol $\dot{t}$ represents the derivative of $\tau$, which is the affine parameter. For null geodesics, i.e., $m=0$, the equations of motion for photons propagating in the spacetime (8) can be obtained by solving the Hamilton-Jacobi equation. Further combined with Eq.(14), the equations of motion for photons can be expressed as the following four first-order differential equations
@@
$R=\left[a^{2}-a\xi+\mathcal{H}\right]^{2}E^{2}-\Delta E^{2}\left(a^{2}-2a\xi+\eta+\xi^{2}\right)=0,$ (19)
@@
$\xi=\frac{a^{2}\Delta^{\prime}-2\Delta\mathcal{H}^{\prime}+\mathcal{H}\Delta^{\prime}}{a\Delta^{\prime}}|_{r=r_{p}},$ (21)
@@
Due to the non-negativity of $\Theta$ in Eq.(15), it immediately presents the condition for the photon region
@@
$e_{3}=-e_{(\phi)}=\left(0,0,0,\frac{-1}{\sqrt{g_{\phi\phi}}}\right),$ (27)
@@
The observer is located at ($r_{O},\theta_{O}$) in the coordinates ($t,r,\theta,\phi$), which corresponding to the point $O$ in Fig.2. The red line with arrow originated from $O$ represents the propagated direction of the light ray. And, the vector $\overrightarrow{OP}$ denotes the tangent vector of the null geodesic at $O$. To further obtain the shadow and image of black hole, we take the $\overrightarrow{OP}$ as a radius and $O$ as the centre to plot a three-dimensional sphere, and place the origin of the tetrad (24) at $O^{\prime}$. The line $O^{\prime}S$ passed through the point $O$, as the diameter of the sphere, is on a line with $e_{1}$. The blue rectangular plane at $O^{\prime}$ is the screen for presenting the shadow and image of black hole. By regarding the point $S$ as the reference point, the vector $\overrightarrow{OP}$ is projected to be the vector ${O^{\prime}P^{\prime}}$ on the screen. To determine the photon’s position from the observer’s perspective, it is useful to introduce celestial coordinates($\alpha,\beta$). For the plane $O^{\prime}PS$ with the green boundary of the sphere, the first celestial coordinate, $\alpha$, is defined as the angle between $OO^{\prime}$ and $OP$. For the brown vertical plane of the sphere, it intersects with the front side of the plane $O^{\prime}PS$ at point $Q$. The other celestial coordinate, $\beta$, is defined as the angle between $OQ$ and $e_{2}$.
For the null geodesic, i.e., $s(\tau)={t(\tau),r(\tau),\theta(\tau),\phi(\tau)}$, its tangent should be a linear combination of $e_{i}$ere $i=0,1,2,3.$, which form can be expressed as
@@
where, the negative sign ensures that the tangent vector is oriented towards the past, and the symbol $\cdot$ denotes the partial derivative with respect to the affine parameter $\tau$. And, we note that the path of the photon does not depend on its energy. Therefore, we set the energy of photon in the camera’s frame, $E$, to be unity, i.e., $E_{camera}=1=|\overrightarrow{OP}|\cdot\chi=-\frac{E}{\sqrt{g_{tt}}}|_{(r_{O},\theta_{O})}$. In addition, the four-momentum of photons in the ZAMO’s frame can be rewrote as $p_{(\mu)}=p_{\nu}e^{\nu}_{(\mu)}$, where $e^{\nu}_{(\mu)}$ is given by Eqs.(24-27). Since the values of the 4-momentum of a photon can be obtained by Eqs.(15-18), we can easy extract the momentum of photons $p_{(\mu)}$. By following the method in [110], the connection between the celestial coordinates and the four-momentum of photons $p_{(\mu)}$ is represented as $\cos\alpha={p^{(1)}}/{p^{(0)}},\tan\beta={p^{(3)}}/{p^{(2)}}.$ Specifically, this two celestial coordinates for the rotating LQG black holes are,
@@
where, $\mathcal{N}=\mathcal{A}-a^{2}\Delta\cos^{2}\theta$, $\mathcal{M}=\mathcal{A}a^{2}\sin^{2}\theta(\mathcal{H}-\mathcal{F}\mathcal{H})^{2}$. With the help of Eq.(31), we introduce the Cartesian coordinates system ($x,y$) on the screen, it yields
@@
In Fig.3, it shows the influence of $\lambda$ on the black hole shadow for different values of $a$. Clearly, one can see that the parameter $\lambda$ always decreases the size of shadow for both $a=0.1$ and $a=0.998$, and seems to have no effects on the deformation of shadow. According to the Fig.1, we know that the acceptable range of values for $\lambda$ is bigger and bigger with the decrease of the rotating parameter $a$. So, when $a=0.1$, the broad range of $\lambda$ leads to substantial variations in the shadow’s size. Conversely, with $a=0.998$, the narrower range of $\lambda$ results in a smaller changes in the shadow size.
(a) $(a)r_{o}=100,\theta_{o}=\pi/2$
(b) $(b)a=0.998,r_{o}=100$
(c) $(c)a=0.998,\theta_{o}=\pi/2$
Figure 4: The shadow versus parameters $a,r_{o},$ and $\theta_{o}$.
@@
Next, we continue to focus on the deviation from the circularity $\delta_{s}$ and the size $R_{s}$ of the shadow cast by the rotating black holes, which defined by Hioki and Maeda [111].
@@
In Fig.5, the five reference points ($\alpha_{t},\beta_{t}$), ($\alpha_{b},\beta_{b}$), ($\alpha_{r},0$), ($\alpha_{p},0$) and ($\alpha_{p^{\prime}},0$) correspond to the top, bottom, rightmost, and leftmost points of the shadow, leftmost point of the reference circle, respectively. The deviation from the circularity ($\delta_{s}$) and the size ($R_{s}$) are defined as
@@
In Tables 1 and 2, we intuitively show the deviation from the circularity $\delta_{s}$ and the size $R_{s}$ of the shadow cast by the non-singular rotating black holes.